OpenStudy (anonymous):

Determine whether a quadratic model exists for each set of values. If so, write the model. f(0)=5, f(2)=3, f(-1)=0

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

you want an answer or a method?

OpenStudy (anonymous):

i just want to know how to get the answer. I was in class and i understood but its just the way this question is set up.

OpenStudy (anonymous):

ok we can go through it carefully if you like start with \[f(x)=ax^2+bx+c\] and use the info you got to find \(a,b, c\)

OpenStudy (anonymous):

So far I have f(0)=a(0)^2+b(0)^2+c=5

OpenStudy (anonymous):

is that how i start?

OpenStudy (anonymous):

ok so now you know \(c=5\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is where i stopped because i dont know where to go

OpenStudy (anonymous):

repeat for the other two, but you will have to use two equations and two unknowns \(a\) and \(b\)

OpenStudy (anonymous):

\[f(x)=ax^2+bx+5\] since we got rid of the \(c\) on the first step \[f(2)=a\times 2^2+b\times 2+5=3\\ 4a+2b+5=3\\ 4a+2b=-2\] is one equation

OpenStudy (anonymous):

repeat for \(x=-1\) and get another equation in \(a\) and \(b\) clear?

OpenStudy (anonymous):

So.. a(-1)^2+b(-1)^2+5=0?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh not right !

OpenStudy (anonymous):

\[a(-1)^2+b(-1)+5=0\] is more like it no square on the second term

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you wrote \[a(-1)^2+b(-1)^2+5=0\] but that is not right, because you had a square on the second term

OpenStudy (anonymous):

it is \[f(x)=ax^2+bx+5\]

OpenStudy (anonymous):

no square on the second term

OpenStudy (anonymous):

Only on the first part because the standard form is ax^2+bx+c?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so the second equation is \[a-b=-5\]

OpenStudy (anonymous):

and the system is \[4a+2b=-2\\ a-b=-5\] which is not too hard to solve, either by elimination or substitution

OpenStudy (anonymous):

btw i hope it is clear that \[a(-1)^2+b(-1)+5=0\]is the same as \[a-b=-5\]

OpenStudy (anonymous):

O

OpenStudy (anonymous):

is this the third part or the second?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

so what happens to the 3 and 0?

OpenStudy (anonymous):

what do you mean? oh i see hold on

OpenStudy (anonymous):

we got this \[a(-1)^2+b(-1)+5=0\] right?

OpenStudy (anonymous):

the f(2)=3 , f(-1)=0 .... what happens to the 3 and 0?

OpenStudy (anonymous):

the zero is in \[a(-1)^2+b(-1)+5=0\] on the right then compute and get \[a-b+5=0\]

OpenStudy (anonymous):

subtracting \(5\) gives \[a-b=-5\]

OpenStudy (anonymous):

\[f(x)=ax^2+bx+5\] \[f(2)=a\times 2^2+b\times 2+5=3\] i.e. \[4a+2b+5=3\] the 3 is on the right subtract 5 get \[4a+2b=-2\]

OpenStudy (anonymous):

im going to look at it for a minute..

OpenStudy (anonymous):

take your time maybe i skipped on step

OpenStudy (anonymous):

\[f(2)=a\times 2^2+b\times 2+5=3\] \[a\times 4+b\times 2+5=3\\ 4a+2b+5=3\\ 4a+2b=-2\]

OpenStudy (anonymous):

OK that makes better sense

OpenStudy (anonymous):

lol it is the old commutative law that confuses every time!

OpenStudy (anonymous):

so what are the equations supposed to look like?

OpenStudy (anonymous):

it was so long ago i forget we just got \[4a+2b=-2\] as one of them right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

aah the other is \[a(-1)^2+b(-1)+5=0\] and i am going to let you figure out why that is the same as \[a-b=-5\]

OpenStudy (anonymous):

system to solve is \[4a+2b=-2\\ a-b=-5\]

OpenStudy (anonymous):

subtract 5 on both sides so it becomes a - and -1 x -1 =1 2 which is just 'a' and do the same for b then im left with a-b=-5... right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can i use substitution?

OpenStudy (anonymous):

can i use substitution?

OpenStudy (anonymous):

wait hold up am i solving for a and b?

OpenStudy (anonymous):

im stuck

OpenStudy (anonymous):

you can use substitution or elimination, your choice

OpenStudy (anonymous):

its 10 at night and im only on the 1st question out of 32....

OpenStudy (anonymous):

they get quicker

OpenStudy (anonymous):

\[4a+2b=-2\\ a-b=-5\] the second equation is the same as \(a=b-5\) so you can substitute in the first one and solve \[4(b-5)+2b=-5\]

OpenStudy (anonymous):

doing it right now ;)

OpenStudy (anonymous):

b=3

OpenStudy (anonymous):

damn i made a typo \[4(b-5)+2b=-2\] is more like it yes \(b=3\)

OpenStudy (anonymous):

and so \(a=-2\) and your answer is \[f(x)=-2x^2+3x+5\] want to check it?

OpenStudy (anonymous):

just wait

OpenStudy (anonymous):

where am i plugging the 3 at

OpenStudy (anonymous):

either of these \[4a+2b=-2\\ a-b=-5\]

OpenStudy (anonymous):

nevermind i just got confused cause i got to0 much on my paper

OpenStudy (anonymous):

\[a-3=-5\iff a=-2\]

OpenStudy (anonymous):

so the question was asking me to determine whether a quadratic model exists, how do i know if it does?

OpenStudy (anonymous):

that is a dumb question unless it was a line or not a function you have three equations and three unknowns you can almost always solve that

OpenStudy (anonymous):

i don't mean you asked a dumb question, i mean the question that you were asked was dumb

OpenStudy (anonymous):

if i said \[f(0)-5, f(1)=4,f(3)=6\] i could make a quadratic, out of it the coefficients might be annoying fractions, not like the simple answer we got, but i could do it

OpenStudy (anonymous):

for the next problem its f(-1)=-4, f(1)=-2, f(2)=-1 and the answer says that there is no model.

OpenStudy (anonymous):

hmm let me see if i can figure out why

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first lets cheat and see that our last answer was correct

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

now that you know how to cheat, lets do it for the one it says can't be done

OpenStudy (anonymous):

k

OpenStudy (anonymous):

oooh doe it is a line!!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so in order to know would i just graph it?

OpenStudy (anonymous):

i guess i wasn't thinking i suppose you could try to make it \[f(x)=ax^2+bx+c\] as before and discover that \(a=0\)

OpenStudy (anonymous):

now if you really have 30 of these to do (that seems like an awful lot) i would cheat my brains out

OpenStudy (anonymous):

when i graph and it turns out as a parabola then a model exists?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

hate being an honors student

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

think of all the honors you'll get not to mention the fact that you avoid sitting in a class full of "i don't get it"s

OpenStudy (anonymous):

yeah lol well just one last question how would i graph the one we spent a long time working out? i know its supposed to become a parabola but when i graph it it becomes funky looking.

OpenStudy (anonymous):

there is a picture in the link i sent use that

OpenStudy (anonymous):

O well thank you! Cant believe you were willing to sit here for an hour+ showing me to work a problem. Again appreciate the help now im going to get these other 31 problems done lol

OpenStudy (anonymous):

@satellite73